OpenStudy (anonymous):

The point (2,y) is a solution to the equation -2x+3y-4=0. What is the value of y? F. -3 G. 0 H. 8/3 J. 4

jimthompson5910 (jim_thompson5910):

(2,y) means x = 2 and y is unknown (we just leave it as y for now)

jimthompson5910 (jim_thompson5910):

Go to the equation -2x+3y-4=0 and plug in x = 2 -2x+3y-4=0 -2(2)+3y-4=0 ... replace every copy of x with 2 -4+3y-4=0 3y - 8 = 0 ... ... ... y = ??

OpenStudy (anonymous):

Is my answer j?

jimthompson5910 (jim_thompson5910):

what do you get when you solve for y in 3y - 8 = 0 ignore the answer choices for now

OpenStudy (anonymous):

Y=-2.6

jimthompson5910 (jim_thompson5910):

write that as a fraction and the answer should be positive (not negative)

OpenStudy (anonymous):

2/6

jimthompson5910 (jim_thompson5910):

that's not correct

OpenStudy (anonymous):

I don't have a calculator to put in a fraction

jimthompson5910 (jim_thompson5910):

3y - 8 = 0 3y - 8+8 = 0+8 ... Add 8 to both sides. 3y = 8 what's the next step?